# Missing Pieces: How Framing Uncertainty Impacts Longitudinal Trust in AI Decision Aids — A Gig Driver Case Study

### v1

@@ Appendix A Survey Questions
A.1. Survey 1 — Formative Questions
@@ A.2. Survey 2 — Pre-Interaction Trust
This item is based on item U2 from the Perceived Understandability questions of (Madsen and Gregor, 2000), “I understand how the system will assist me with decisions I have to make.” As the constraint page is intended to encapsulate the user’s decision-making process, we consider the generation of the recommended schedule to be how the tool assists the user with their decisions.
@@
This item is based on item R4 from the Perceived Reliability questions of (Madsen and Gregor, 2000), “I can rely on the system to function properly.” We consider the tool to be properly functioning if its recommendations account for the user’s goals and preferences.
@@
This item is based on item T3 from the Perceived Technical Competence questions of (Madsen and Gregor, 2000), “The advice the system produces is as good as that which a highly competent person could produce.” In our context, the tool’s advice is its recommended schedule of driving times, and to our participants a competent individual would be an experienced driver.
@@
This item is based on item F1 from the Faith questions of (Madsen and Gregor, 2000), “I believe advice from the system even when I don’t know for certain that it is correct.” Again, in our context, the tool’s advice is its recommended schedule of driving times. Since we cannot apply a clear notion of correctness to continuous estimates of earnings, we reworded this question to focus on alignment with the user’s objectives.
@@
This item is based on item P4 from the Personal Attachment questions of (Madsen and Gregor, 2000), “I like using the system for decision making.” We reworded the question to better assess participants’ level of intended future reliance on the tool.
A.3. Survey 3 — Daily Trust
@@
This item is based on item U4 of the Perceived Understandability questions of (Madsen and Gregor, 2000), “It is easy to follow what the system does.” Instead of asking the user about the tool’s operation generally, we focused the question on the interpretability of its recommended schedule for that day.
@@
This item is based on item R1 of the Perceived Reliability questions of (Madsen and Gregor, 2000), “The system always provides the advice I require to make my decision.” Again, in our context, the tool’s advice is its recommended schedule. We focused the question on the user’s decisions for that particular day.
@@
This item is based on item F2 from the Faith questions of (Madsen and Gregor, 2000), “When I am uncertain about a decision I believe the system rather than myself.” We focused the question on the user’s decisions for that particular day, and reworded “believe” to “follow” to assess compliance more clearly.
@@
A.4. Survey 4 — Post-Interaction Trust
@@
This item is based on item 12 of (Jian et al., 2000)’s instrument, “I am familiar with the system.” We reworded the question in light of the fact that users did not have any existing experience with using the tool before the study.
@@
This item is based on item 11 of (Jian et al., 2000)’s instrument, “I can trust the system.” We focused the scope of this question on the output of the tool, the recommended schedule, rather than the tool as a whole.
@@
This item is based on item 9 of (Jian et al., 2000)’s instrument, “The system is dependable.” Again, we focused the scope of this question on the output of the tool, the recommended schedule.
@@
This item is an original question that prompts the participant to consider their interactions with other types of recommendation systems. It assesses the extent to which participants would appreciate increased error visibility through ranged estimates in such systems.
@@
This item is based on item 1 of (Jian et al., 2000)’s instrument, “The system is deceptive.” In addition to focusing the scope of this question on the output of the tool, we also reworded “deceptive” to “misleading” to capture the broader possibility of the tool being perceived as unintentionally providing incorrect information.
@@
This item is based on item 5 of (Jian et al., 2000)’s instrument, “The system’s actions will have a harmful or injurious outcome.” In our context, the outcome for the user is their earnings from driving while following the recommended schedule. We reworded the question to assess the outcome retrospectively.
@@
Appendix B Interview Scripts
B.1. Pilot
B.1.1. Formative Questions
@@ B.1.2. Evaluative Questions
B.2. Longitudinal Trials
@@ B.2.1. Formative Questions
B.2.2. Feedback on Constraint Design
@@
Could you see yourself spending more time interacting with the tool than you did (e.g., to enter more details)? Why or why not?
B.2.3. Feedback on Schedules
@@ B.2.4. Feedback on Estimates
If participants were in Conditions 1 or 2 Would you have preferred to see a range for how much you could earn?
If participants were in Conditions 3 or 4 Would you have preferred to see a single number for how much you could earn?
@@ B.2.5. Overall Thoughts
Appendix C Confounding Variable Analysis
In this section, we detail the models we used to analyse whether several confounding variables could have impacted our measurements. We use the variable notation introduced in Section 6.2.1; for a participant’s selected constraint variant; for the participant’s city; and for the participant’s level of motivation, on a scale of 1–5.
C.1. Constraint Variants
To ascertain that trust and reliance measures did not significantly vary between different variants of constraints (maximum hours or minimum earnings/daily or weekly), we encoded these variants of the constraints as a categorical variable , and added it as linear terms to the best-fitting models for each measurement:
Survey 2 Trust (Table S1): OLS model from RQ3
Survey 3 Trust (Table S2): LMM from RQ1
Survey 3 Reliance (Table S3): LMM modified from RQ2
Survey 4 Trust (Table S4): OLS model from RQ1
These revised models generally led to worse AIC values (with models including interactions of being worse still), and most of the coefficients for the constraint variants were not significant. As noted in Section 6.2.1, one exception was the max/day group, who exhibited significantly higher reliance for Survey 3. The AIC was slightly improved by adding the constraints to this model. Fitted coefficients for each model are shown in the tables listed above.
Coefficient | |||
Intercept | 3.1846 | 0.376 | 0.000*** |
0.4304 | 0.423 | 0.315 | |
0.0459 | 0.338 | 0.893 | |
0.0317 | 0.433 | 0.942 | |
0.3341 | 0.383 | 0.387 | |
0.9385 | 0.403 | 0.024* | |
0.1244 | 0.409 | 0.762 |
Table of ordinary least squares model coefficients for pre-interaction Survey 2 trust, with constraint added. Significant coefficients are the intercept (3.1846) and Condition 3 (0.9385), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 2.992 | 0.271 | 0.000*** |
0.216 | 0.268 | 0.421 | |
0.303 | 0.359 | 0.399 | |
0.540 | 0.403 | 0.180 | |
0.129 | 0.044 | 0.003** | |
0.005 | 0.002 | 0.003** | |
-0.001 | 0.000 | 0.008** | |
0.517 | 0.232 | 0.026* | |
-0.116 | 0.050 | 0.019* | |
Random effect | 0.499 | 0.246 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust, with constraint added. Significant coefficients are the intercept (2.992), (0.129), (0.005), the interaction of and (-0.001), (0.517), and the interaction of and (-0.116), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 0.431 | 0.088 | 0.000*** |
0.133 | 0.101 | 0.186 | |
0.322 | 0.134 | 0.016* | |
-0.144 | 0.153 | 0.346 | |
0.230 | 0.073 | 0.002** | |
0.058 | 0.066 | 0.383 | |
Random effect | 0.059 | 0.073 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance, with constraint added. Significant coefficients are the intercept (0.431), (0.23), and = max/day (0.322).
Coefficient | |||
Intercept | 3.539 | 0.345 | 0.000*** |
-0.021 | 0.256 | 0.934 | |
0.267 | 0.323 | 0.416 | |
-1.139 | 0.732 | 0.131 | |
0.274 | 0.349 | 0.439 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust, with constraint added. The intercept (3.539) is significant, but the coefficients for (0.274) and are not.
C.2. Cities
To ascertain that trust and reliance measures did not significantly vary based on the city in which participants were driving (Los Angeles, New York, Chicago, or Houston), we similarly added a categorical variable for the city to the best-fitting models:
Survey 2 Trust (Table S5): OLS model
Survey 3 Trust (Table S6): LMM
Survey 3 Reliance (Table S7): LMM
Survey 4 Trust (Table S8): OLS model
These revised models all led to worse AIC values, and none of the coefficients for the cities were significant.
Coefficient | |||
Intercept | 3.0656 | 0.467 | 0.000*** |
0.4695 | 0.497 | 0.350 | |
0.2597 | 0.417 | 0.537 | |
0.3459 | 0.494 | 0.488 | |
0.2646 | 0.394 | 0.505 | |
0.8850 | 0.430 | 0.045* | |
0.0353 | 0.405 | 0.931 |
Table of ordinary least squares model coefficients for pre-interaction Survey 2 trust, with city added. Significant coefficients are the intercept (3.0656) and Condition 3 (0.885), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 3.121 | 0.371 | 0.000*** |
0.174 | 0.440 | 0.693 | |
0.015 | 0.351 | 0.966 | |
0.053 | 0.402 | 0.895 | |
0.132 | 0.044 | 0.003** | |
0.004 | 0.002 | 0.004** | |
-0.001 | 0.000 | 0.011* | |
0.533 | 0.233 | 0.022* | |
-0.120 | 0.050 | 0.016* | |
Random effect | 0.525 | 0.257 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust, with city added. Significant coefficients are the intercept (3.121), (0.132), (0.004), the interaction of and (-0.001), (0.533), and the interaction of and (-0.12), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 0.539 | 0.137 | 0.000*** |
0.002 | 0.179 | 0.990 | |
-0.033 | 0.143 | 0.818 | |
-0.023 | 0.163 | 0.888 | |
0.213 | 0.073 | 0.004** | |
0.067 | 0.066 | 0.308 | |
Random effect | 0.077 | 0.090 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance, with city added. Significant coefficients are the intercept (0.539) and (0.213), but not .
Coefficient | |||
Intercept | 3.4142 | 0.389 | 0.000*** |
0.1228 | 0.410 | 0.767 | |
0.0096 | 0.321 | 0.976 | |
-0.3000 | 0.388 | 0.445 | |
0.4629 | 0.331 | 0.173 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust, with city added. The intercept (3.4142) is significant, but the coefficients for (0.4629) and are not.
C.3. Motivation
To ascertain that trust and reliance measures did not significantly vary with participants’ level of motivation, we needed to first select a survey item that would allow us to assess this subjective dimension independent of the participants’ perception of the tool. For this reason, we selected Question 5 of the formative Survey 1 (Section A.1), “When I drive, it’s important to me that I make some minimum amount of money.” We encoded the Likert responses to this question as an interval variable and added it as linear terms to the best-fitting models for each measurement:
Survey 2 Trust (Table S9): OLS model from RQ3
Survey 3 Trust (Table S10): LMM from RQ1
Survey 3 Reliance (Table S11): LMM modified from RQ2
Survey 4 Trust (Table S12): OLS model from RQ1
These revised models again generally led to worse AIC values, and most of the coefficients for the motivation variable were not significant. As noted in Section 6.2.1, one exception was the pre-interaction trust score, which had a significant positive correlation with the motivation variable. The AIC was improved by adding the motivation variable to this model. However, no such effects were evident for the daily and post-interaction measures. This suggests that the initial effects of the participants’ motivation on trust did not persist over repeated interactions. Fitted coefficients for each model are shown in the tables listed above.
Coefficient | |||
Intercept | 2.1844 | 0.475 | 0.000*** |
0.3001 | 0.105 | 0.007** | |
0.0725 | 0.352 | 0.838 | |
0.7487 | 0.359 | 0.043* | |
-0.0430 | 0.358 | 0.905 |
Table of ordinary least squares model coefficients for pre-interaction Survey 2 trust, with motivation added. Significant coefficients are the intercept (2.1844), Condition 3 (0.7487), and (0.3001).
Coefficient | |||
Fixed effects | |||
Intercept | 2.990 | 0.570 | 0.000*** |
0.046 | 0.119 | 0.699 | |
0.126 | 0.044 | 0.004** | |
0.002 | 0.001 | 0.001*** | |
-0.000 | 0.000 | 0.004** | |
0.530 | 0.230 | 0.021* | |
-0.116 | 0.049 | 0.018* | |
Random effect | 0.530 | 0.261 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 trust, with motivation added. Significant coefficients are the intercept (2.99), (0.126), (0.002), the interaction of and (-0.000), (0.53), and the interaction of and (-0.116), but not .
Coefficient | |||
Fixed effects | |||
Intercept | 0.383 | 0.234 | 0.102 |
0.030 | 0.050 | 0.550 | |
0.211 | 0.071 | 0.003** | |
0.074 | 0.064 | 0.250 | |
Random effect | 0.081 | 0.094 |
Table of linear mixed model coefficients (intercept, fixed effects, and random effects) for daily Survey 3 reliance, with motivation added. The fixed effect for (0.211) is significant, but not .
Coefficient | |||
Intercept | 3.9645 | 0.676 | 0.000*** |
-0.1259 | 0.127 | 0.330 | |
0.4258 | 0.317 | 0.189 |
Table of ordinary least squares model coefficients for post-interaction Survey 4 trust, with motivation added. The intercept (3.9645) is significant, but the coefficients for (0.4258) and are not.
Appendix D Final Tool Design
The following figures show screenshots of the web-based tool used for the longitudinal trials (Section 6), showing the constraints and recommended schedule of interview participant F1.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the constraint page is shown in a grey iframe. Various inputs for constraints are set to values entered by interview participant F1. In response to “Would you like to specify a maximum number of hours that you would like to drive, or would you prefer to specify a minimum amount of money that you would like to earn?”, F1 selected the latter. In response to “Would you like to set these constraints for each day individually, or for the entire week?”, F1 selected the latter. In response to “What is the minimum amount of money you would like to earn for the entire week?”, F1 entered $1 400. In response to “When are you available during the week?”, F1 selected time slots (as indicated by blue cells in a tabular schedule) corresponding to 3 am to 7 am, 8 am to 4 pm, and 10 pm to 12 am on every day of the week.
D.1. Schedule Page Conditions
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 1 for interview participant F1 is shown in a grey iframe. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn a weekly average of”, followed by an estimate of $1 412 below. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 2 for interview participant F1 is shown in a grey iframe. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn”. Unlike Condition 1, this is followed by a table that reads, “An average of”, alongside estimates for each day: $210 for Monday, $194 for Tuesday, $175 for Wednesday, $173 for Thursday, $194 for Friday, $226 for Saturday, and $237 for Sunday. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 3 for interview participant F1 is shown in a grey iframe. The top of the schedule page reads, “Based on historical data, it is estimated that you will earn”. Unlike Conditions 1 and 2, this is followed by a table that reads, “On an average week: $1 412”, “On a bad week: $391”, and “On a good week: $3 032”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Unlike Conditions 1 and 2, this is also followed by “Low”, with the 10th percentile of earnings, and “High”, with the 90th percentile of earnings. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.
Screenshot of a webpage titled “Tool Interaction”. Below a set of instructions, the schedule page in Condition 4 for interview participant F1 is shown in a grey iframe. Unlike Conditions 1, 2, and 3, the top of the schedule page reads, “In general, every driver may encounter different conditions while they are driving, which leads to differences in earnings from individual to individual and from one day to the next. On average, based on historical data, a driver following this schedule will earn”. Also unlike Conditions 1 and 2, this is followed by a table that reads, “On an average week: $1 412”, “On a bad week: $391”, and “On a good week: $3 032”. Further below, a tabular schedule is shown for each day of the week. Each cell reads, “Average”, followed by the estimated hourly earning in bold. Unlike Conditions 1 and 2, this is also followed by “Low”, with the 10th percentile of earnings, and “High”, with the 90th percentile of earnings. Recommended time slots are indicated by blue cells, non-recommended time slots during which F1 is still available are indicated by dark text, and time slots during which F1 is not available are indicated by light text.

### v2

@@ Appendix A Survey Questions
A.1. Survey 1 — Intake Survey
@@ A.2. Survey 2 — Pre-Survey
This item is based on item U2 from the Perceived Understandability questions of the HCT (Madsen and Gregor, 2000), “I understand how the system will assist me with decisions I have to make.” As the constraint page is intended to encapsulate the user’s decision-making process, we consider the generation of the recommended schedule to be how the tool assists the user with their decisions.
@@
This item is based on item R4 from the Perceived Reliability questions of the HCT (Madsen and Gregor, 2000), “I can rely on the system to function properly.” We consider the tool to be properly functioning if its recommendations account for the user’s goals and preferences.
@@
This item is based on item T3 from the Perceived Technical Competence questions of the HCT (Madsen and Gregor, 2000), “The advice the system produces is as good as that which a highly competent person could produce.” Our tool’s advice is its recommended schedule, and to our participants a competent individual would be an experienced driver.
@@
This item is based on item F1 from the Faith questions of the HCT (Madsen and Gregor, 2000), “I believe advice from the system even when I don’t know for certain that it is correct.” Again, our tool’s advice is its recommended schedule of driving times. Since we cannot apply a clear notion of correctness to continuous estimates of earnings, we reworded this question to focus on alignment with the user’s objectives.
@@
This item is based on item P4 from the Personal Attachment questions of the HCT (Madsen and Gregor, 2000), “I like using the system for decision making.” We reworded it to better assess participants’ level of intended future reliance on the tool.
A.3. Survey 3 — End-of-Day Survey
@@
This item is based on item U4 of the Perceived Understandability questions of the HCT (Madsen and Gregor, 2000), “It is easy to follow what the system does.” Instead of asking the user about the tool’s operation generally, we focused the question on the interpretability of its recommended schedule for that day.
@@
This item is based on item R1 of the Perceived Reliability questions of the HCT (Madsen and Gregor, 2000), “The system always provides the advice I require to make my decision.” Again, our tool’s advice is its recommended schedule. We focused the question on the user’s decisions for that particular day.
@@
This item is based on item F2 from the Faith questions of the HCT (Madsen and Gregor, 2000), “When I am uncertain about a decision I believe the system rather than myself.” We focused the question on the user’s decisions for that particular day, and reworded “believe” to “follow” to assess compliance more clearly.
@@
A.4. Survey 4 — Post-Survey
@@
This item is based on item 12 of the TiA (Jian et al., 2000), “I am familiar with the system.” We reworded the question in light of the fact that users did not have any existing experience with using the tool before the study.
@@
This item is based on item 11 of the TiA (Jian et al., 2000), “I can trust the system.” We focused the scope of this question on the output of the tool, the recommended schedule, rather than the tool as a whole.
@@
This item is based on item 9 of the TiA (Jian et al., 2000), “The system is dependable.” Again, we focused the scope of this question on the output of the tool, the recommended schedule.
@@
This item is an original question that prompts the participant to consider their interactions with other types of recommendation systems. It assesses the extent to which participants would appreciate increased exposure of uncertainty through ranged estimates in such systems.
@@
This item is based on item 1 of the TiA (Jian et al., 2000), “The system is deceptive.” In addition to focusing the scope of this question on the output of the tool, we also reworded “deceptive” to “misleading” to capture the broader possibility of the tool being perceived as unintentionally providing incorrect information.
@@
This item is based on item 5 of the TiA (Jian et al., 2000), “The system’s actions will have a harmful or injurious outcome.” In our context, the outcome for the user is their earnings from driving while following the recommended schedule. We reworded the question to assess the outcome retrospectively.
@@
Appendix B Pilot Interviews
To assess the utility of the AI-based schedule recommendation tool design introduced in Section 3, we created a prototype based on it. We then conducted a series of interviews to understand gig drivers’ needs and how well the prototype aligned with them. The interview methodology was approved by our Institutional Review Board (IRB).
B.1. Tool Design
The constraint page’s design (Figure 7) was largely similar to the final design. However, the prototype had a monotone colour scheme and had instructions that were worded less clearly, which were improved based on feedback from the pilot (Section B.5). Additionally, due to technical limitations, dropdowns were used in place of text boxes. However, we consider the impact of these limitations to be minor since the interviewer controlled the page.
The schedule page’s design (Figure 8) was kept static for the pilot to gather more uniform feedback from participants. Like the constraint page, the prototype differed from the final design in its colour scheme and clarity of wording. The prototype was most similar to Condition (R) (Section 3.3) in that it displayed a range consisting of mean (“On an average week…”), pessimistic (“On a bad week…”), and optimistic (“On a good week…”) weekly earnings above the tabular schedule. Unlike the final design of Condition (R), however, ranges were not shown for hourly estimates; this was added based on feedback from the pilot (Section B.5).
B.2. Methodology
Participants began by completing a web-based consent form and a demographics survey that collected their age, gender, and education level. This form was separate from the one for the longitudinal user study (Section 4). After completing the web form, participants were invited via email to complete 20–30-minute audio-recorded Zoom interviews conducted by a single author. Participants were compensated with a $10 Amazon gift card.
In the first 5–10 minutes, the interview focused on formatively understanding drivers’ needs and motivations. The questions asked in this portion of the interview were the same as the Intake Survey of the longitudinal user studies (Section 4.2). In the last 15–20 minutes, the interview focused on evaluatively understanding how well the tool met drivers’ needs. To ensure a consistent experience, the Figma prototype was opened in the interviewer’s browser and shown to participants in a screensharing session. First, on the constraint page, the participant was asked to work with the interviewer to interact with the page, entering the constraints as if they were using the tool for their actual planning. Then, on the schedule page, the participant was shown a schedule with mocked earnings estimates. Finally, the participant was asked about their overall opinions of the tool. The full interview script is shown in Section B.3.
To analyse these interviews, we used the same methodology as the longitudinal interviews (Section 4.3).
B.3. Interview Script
B.3.1. Formative Questions
@@ B.3.2. Evaluative Questions
B.4. Participants
As with the longitudinal user studies (Section 4.1), participants were recruited from the user base of Gridwise in July 2023. Gridwise distributed recruitment messages to 500 users, but otherwise did not interact with participants. Recipients were sampled from users in the United States who had completed at least one gig in a platform linked to the Gridwise app over the week preceding recruitment. We recruited 4 interview participants:
B.5. Results
The four pilot interview participants reported a diversity of motivations and routines for driving. While all four participants had specific earning goals, P1, P2, and P4 considered their goals to be important and valued the time flexibility of gig work, whereas P3 was more motivated by the opportunity for human interaction. P1, P2 and P3 had typical times that they drive at; however, P1, P2, and P4 also adjusted their schedules based on demand. All four participants had encountered difficulties in planning due to the unpredictability of demand and/or supply (with P1, P2, and P3 feeling that gig platforms provide insufficient information), and indicated that they would find schedule recommendations to be useful.
All four participants found the initial design of the tool to be generally understandable, and felt that it would be useful for drivers in planning their activity. P1 and P2 liked the fact that the tool presents information to them in a way that reduces the need for guesswork while driving. In particular, P1 suggested that the tool would help mitigate a catch-22: it is not possible to view gig demand information in DoorDash without exiting their Dash (scheduled work period), but doing so seemingly deprioritises them.
On the constraint page, P1 and P4 indicated that the questions aligned well with their goals. P1 and P3 suggested that they would not set the constraints to perfectly align with their routines, so as to receive more information from the tool. On the schedule page, all four participants liked the estimated hourly earnings, with P1, P2, and P3 indicating that they would be helpful in deciding whether or not to work at particular times of day. Yet, P2, P3, and P4 acknowledged that the estimates would only be guesses. P1 and P4 also liked the range of weekly earnings, but P3 felt it assumed they would follow the recommended schedule perfectly. P2 and P3 noted that ranges for hourly earnings would be useful to display.
More granular constraints. P1, P3, and P4 all suggested ways to limit the scope of the historical gigs used to estimate their earnings. P3 wanted the tool to clarify that the historical data was limited to the city they drive in, and also how recent the data was. P1 and P3 wanted to limit the maximum distance of the historical gigs from their starting point. P4, who works for less popular delivery platforms, wanted to select which platforms the historical gigs came from, and indicated that this would improve their perceived control.
Feedback on performance. P1 and P2 both wanted a way to compare the tool’s estimates with their actual earnings. Regardless of how the estimates compared to reality, P1 suggested that this feedback would be motivating; both P1 and P2 remarked that they have gamified their gig-driving experiences to compare against either themselves or others. P2 also wanted to compare the estimates with their expenses, and P3 wanted a way to view the overall supply of drivers.
Appendix C User Study Interview Script
@@ C.1. Formative Questions
C.2. Feedback on Constraint Design
@@
Could you see yourself spending more time interacting with the tool than you did (e.g. to enter more details)? Why or why not?
C.3. Feedback on Schedules
@@ C.4. Feedback on Estimates
[If participants were in Conditions (B) or (D)] Would you have preferred to see a range for how much you could earn?
[If participants were in Conditions (R) or (RH)] Would you have preferred to see a single number for how much you could earn?
@@ C.5. Overall Thoughts
Appendix D Full Quantitative Results
D.1. Daily Trust and Reliance Statistics
In Table 3, we report the means, standard errors, and 95% confidence intervals of the daily trust and reliance scores plotted in Figures 4 and 6. We show both the overall statistics as well as the statistics for each condition. In Figure 6, we show visualisations for daily reliance scores omitted from the main text.
D.2. Trust and Reliance Score LMMs
In Table 1 and Table 2, we show the fitted coefficients for the trust and reliance score LMMs discussed in Section 5.2 and Section 5.3.2.
Factor | Without Condition | With Condition | ||||
SE | SE | |||||
Intercept | 0.00337** | 0.00068*** | ||||
pre_trust_score | 0.00029*** | 0.00256** | ||||
day | ¡0.00001*** | 0.25896 | ||||
estimate_accurate | 0.01357* | 0.02010* | ||||
day:estimate_accurate | 0.00126** | 0.00168** | ||||
condition(D) | 0.15139 | |||||
condition(R) | 0.64038 | |||||
condition(RH) | 0.15755 | |||||
day:condition(D) | 0.03548* | |||||
day:condition(R) | 0.06976† | |||||
day:condition(RH) | 0.14971 | |||||
Random intercept SD |
Factor | Without Condition | With Condition | ||||
SE | SE | |||||
Intercept | 0.95494 | 0.79607 | ||||
pre_trust_score | 0.23316 | 0.27397 | ||||
day | 0.01444* | 0.02280* | ||||
estimate_accurate | 0.50517 | 0.50469 | ||||
trust_score | 0.24285 | 0.38470 | ||||
day:estimate_accurate | 0.81042 | 0.86786 | ||||
day:trust_score | 0.01640* | 0.07127† | ||||
condition(D) | 0.92461 | |||||
condition(R) | 0.63782 | |||||
condition(RH) | 0.63414 | |||||
day:condition(D) | 0.35761 | |||||
day:condition(R) | 0.05958† | |||||
day:condition(RH) | 0.50832 | |||||
Random intercept SD |
D.3. Post-Interaction Trust
We fitted an ordinary least squares (OLS) models for the post_trust_score using statsmodels. This model included the condition along with all previous trust (pre_trust_score and daily trust_score) and reliance measurements.
For Conditions (B)/(D)/(R)/(RH), the mean post-interaction trust scores were 4.000, 3.720, 3.857, and 3.618. None of these conditions were significantly different from each other, and the coefficients for previous trust and reliance scores were not statistically significant either. The questions we adapted from the TiA asked participants to consider the entire duration of their interaction with the schedule recommendation tool. As a result, this broad, retrospective reflection may have failed to capture more nuanced longitudinal changes in trust and reliance like those we described in Section 5.
Day | Condition | Trust | Reliance | ||||
SE | 95% CI | SE | 95% CI | ||||
1 | Overall | 3.353 | 0.196 | (2.872, 3.834) | 0.176 | 0.107 | (-0.100, 0.453) |
(B) | 3.556 | 0.444 | (2.468, 4.643) | 0.167 | 0.307 | (-0.623, 0.957) | |
(D) | 3.033 | 0.390 | (2.080, 3.986) | 0.200 | 0.200 | (-0.314, 0.714) | |
(R) | 3.810 | 0.348 | (2.959, 4.660) | 0.143 | 0.261 | (-0.528, 0.813) | |
(RH) | 3.242 | 0.379 | (2.315, 4.170) | 0.182 | 0.182 | (-0.286, 0.649) | |
2 | Overall | 3.578 | 0.167 | (3.169, 3.988) | 0.235 | 0.095 | (-0.009, 0.479) |
(B) | 3.722 | 0.416 | (2.703, 4.741) | 0.333 | 0.333 | (-0.524, 1.190) | |
(D) | 3.567 | 0.205 | (3.065, 4.069) | 0.200 | 0.200 | (-0.314, 0.714) | |
(R) | 3.857 | 0.397 | (2.885, 4.830) | 0.429 | 0.202 | (-0.091, 0.948) | |
(RH) | 3.333 | 0.362 | (2.447, 4.220) | 0.091 | 0.091 | (-0.143, 0.325) | |
3 | Overall | 3.500 | 0.168 | (3.089, 3.911) | 0.118 | 0.110 | (-0.165, 0.400) |
(B) | 3.833 | 0.331 | (3.025, 4.642) | -0.167 | 0.307 | (-0.957, 0.623) | |
(D) | 3.400 | 0.364 | (2.508, 4.292) | 0.100 | 0.180 | (-0.361, 0.561) | |
(R) | 4.000 | 0.309 | (3.245, 4.755) | 0.571 | 0.202 | ( 0.052, 1.091) | |
(RH) | 3.091 | 0.270 | (2.430, 3.752) | 0.000 | 0.191 | (-0.490, 0.490) | |
4 | Overall | 3.755 | 0.148 | (3.392, 4.118) | -0.029 | 0.123 | (-0.346, 0.287) |
(B) | 3.778 | 0.306 | (3.028, 4.527) | -0.167 | 0.307 | (-0.957, 0.623) | |
(D) | 3.667 | 0.157 | (3.282, 4.051) | -0.200 | 0.200 | (-0.714, 0.314) | |
(R) | 4.238 | 0.347 | (3.390, 5.086) | 0.286 | 0.286 | (-0.449, 1.020) | |
(RH) | 3.515 | 0.334 | (2.697, 4.333) | 0.000 | 0.234 | (-0.600, 0.600) | |
5 | Overall | 3.588 | 0.159 | (3.199, 3.978) | 0.147 | 0.120 | (-0.162, 0.457) |
(B) | 3.278 | 0.475 | (2.116, 4.439) | -0.167 | 0.307 | (-0.957, 0.623) | |
(D) | 3.733 | 0.257 | (3.104, 4.363) | 0.100 | 0.180 | (-0.361, 0.561) | |
(R) | 4.143 | 0.290 | (3.434, 4.852) | 0.429 | 0.297 | (-0.336, 1.193) | |
(RH) | 3.273 | 0.273 | (2.605, 3.940) | 0.182 | 0.226 | (-0.400, 0.764) | |
6 | Overall | 3.667 | 0.163 | (3.268, 4.065) | 0.088 | 0.122 | (-0.226, 0.402) |
(B) | 3.333 | 0.487 | (2.142, 4.525) | -0.333 | 0.333 | (-1.190, 0.524) | |
(D) | 3.733 | 0.276 | (3.059, 4.408) | 0.200 | 0.200 | (-0.314, 0.714) | |
(R) | 4.143 | 0.340 | (3.311, 4.975) | 0.714 | 0.184 | ( 0.240, 1.188) | |
(RH) | 3.485 | 0.275 | (2.813, 4.157) | -0.182 | 0.182 | (-0.649, 0.286) | |
7 | Overall | 3.843 | 0.149 | (3.478, 4.208) | N/A | N/A | N/A |
(B) | 3.889 | 0.351 | (3.029, 4.749) | N/A | N/A | N/A | |
(D) | 3.733 | 0.247 | (3.128, 4.339) | N/A | N/A | N/A | |
(R) | 4.381 | 0.286 | (3.682, 5.080) | N/A | N/A | N/A | |
(RH) | 3.576 | 0.292 | (2.862, 4.289) | N/A | N/A | N/A |